OpenStudy (anonymous):

Calculate the Cal(kcal) in 1 cup of whole milk: 12 g of carbohydrate, 9 g of fat, and 9 g of protein.

OpenStudy (anonymous):

From http://www.nutristrategy.com/nutrition/calories.htm: "Caloric content of fat, protein, carbohydrates and alcohol. Calories are needed to provide energy so the body functions properly. The number of calories in a food depends on the amount of energy the food provides. The number of calories a person needs depends on age, height, weight, gender, and activity level. People who consume more calories than they burn off in normal daily activity or during exercise are more likely to be overweight. Fat: 1 gram = 9 calories Protein: 1 gram = 4 calories Carbohydrates: 1 gram = 4 calories Alcohol: 1 gram = 7 calories" So, (12 g carbohydrates)*4 Cal=48 Cal; (9 g fat) *9 Cal=81 Cal; (9 g protein)*4 Cal=36 Cal 48 Cal + 81 Cal + 36 Cal = 165 Cal in one glass of milk